OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 0 −3 3 2

OpenStudy (anonymous):

@Mertsj

OpenStudy (espex):

You need to take your two equations and eliminate one of the variables, solve for the remaining, then substitute into the second equation.

OpenStudy (espex):

4x + 2y = 12 x – y = 3 => *2 gives you 2x – 2y = 6 add the new equation 2 from equation 1, 4x + 2y = 12 2x – 2y = 6 ------------ 6x + 0 = 18 => solve for x

OpenStudy (amistre64):

if you eliminate the correct variable, then no substitution is needed to find the value of x :)

OpenStudy (anonymous):

x =3 @eSpeX ?

OpenStudy (espex):

Yep, good job. :)

OpenStudy (anonymous):

Thanks!

OpenStudy (espex):

You're welcome.